Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
A rectangular window is 10 inches longer than it is wide. If the area of the window is 1200 square inches, find the length and the width.

Answer by ankor@dixie-net.com(22740)

Solve using the five-step problem-solving process.
Show all steps necessary to arrive at your solution.
A rectangular window is 10 inches longer than it is wide.
If the area of the window is 1200 square inches, find the length and the width.
:
Let x = the width of the window
then
(x+10) = the length
:
Area: L * W, therefore
x(x+10) = 1200
x^2 + 10x = 1200
x^2 + 10x - 1200 = 0
Factor
(x + 40)(x-30) = 0
Positive solution
x = 30 inches is the width
:
The width is 30" and the length is 10" more, 40"
:
Check by finding the area: 30 * 40 = 1200 sq/in